# boxedasideAside

### v1

@@ 3.1 The boundary of $\mathcal{B}_{E,g}$
Let $\Delta=\overline{\mathcal{M}}_{g}-\mathcal{M}_{g}=\Delta_{0}\cup\Delta_{1}\cup\ldots\cup\Delta_{\lfloor g/2\rfloor}$, denote the boundary of ¯$\overline{\mathcal{M}}_{g}$, with $\Delta_{i}$ its irreducible components. The generic point of $\Delta_{0}$ corresponds to an irreducible curve $C_{0}$ with an ordinary double point, whose normalization $\tilde{C}_{0}$ is a genus-$(g-1)$ curve, while the generic point of $\Delta_{i}$ corresponds to a reducible curve with components of genus $i$ and $g-i$ for any $1\leq i\leq\left\lfloor\frac{g}{2}\right\rfloor$.
By abuse of notation, we use $\Xi_{g_{1},g_{2}-1}$, $\Delta_{g_{1},g_{2}}$ for $g_{1},g_{2}\geq 0$ with $g_{1}+g_{2}=g$ to denote the image of $\Xi_{g_{1},g_{2}-1}$, $\Delta_{g_{1},g_{2}}$ under the forgetful morphism $\overline{M}_{E,g}^{red}\to\overline{\mathcal{M}}_{g}$. Recall that all the boundary components consist of isomorphism classes of curves in the closure of the locus of double covers of the configuration
@@
with exactly $22g-$ 2 distinct smooth branch points and that curves corresponding to elements in the loci $\Xi_{g_{1},g_{2}-1}$ (resp. $\Delta_{g_{1},g_{2}}$) are unramified (resp. ramified) above the node. We see that ¯$\overline{\mathcal{B}}_{E,g}-\mathcal{B}_{E,g}$ consists of the union of the loci $\Xi_{g_{1},g_{2}-1}$ and $\Delta_{g_{1},g_{2}}$. The dimensions of these loci equal $42g-$ 4 by Proposition 2.8. Let $\delta_{1,g-2}\subseteq\Xi_{1,g-2}$ denote the union of the components of $\Xi_{1,g-2}$ whose generic points are each represented by a curve of compact type, and denote $\xi_{1,g-2}=\Xi_{1,g-2}-\delta_{1,g-2}$. We have the following relations:
@@
Additionally, we have $\delta_{1,2}\subseteq\Delta_{2}$ for $g=4$.
Example 3.3 (Description of the boundary ¯$\overline{\mathcal{B}}_{E,2}-\mathcal{B}_{E,2}$).
Let $E=(E,O_{E})$ be an arbitrary elliptic curve as above, and let $f_{E}$ be its $p$-rank. In Example 2.5, we describe the boundary of $\overline{M}_{E,2}-{M}_{E,2}$ of $M_{E,2}$, which we use to describe its image in ¯$\overline{\mathcal{M}}_{2}$ here. The locus ¯$\overline{\mathcal{B}}_{E,2}-\mathcal{B}_{E,2}$ is 00-dimensional and it consists of isomorphism classes of curves that are double covers of
@@
We note that there are exactly $3$ curves of the first kind and a unique curve of the second kind, for any $p>2$. Namely, the étale double covers of $E$ correspond to the choices of $2$-torsion points on $E$. Since #$\#E[2](k)=4$, there are $3$ non-trivial $2$-torsion points on $E$ defining different étale double covers $E^{\prime}\to E$, while $O_{E}$ is the fourth $2$-torsion point, and it defines a trivial double cover $(E,O_{E})\sqcup(E,O_{E})\to(E,O_{E})$. Since $f_{E^{\prime}}=f_{E}$, note that $f_{D_{1}}=f_{E}+1$, while $f_{D_{2}}=2\cdot f_{E}$.
3.2 Computing the dimensions of the $p$-rank strata of ¯$\overline{\mathcal{B}}_{E,g}$
Let $V_{f}(\overline{\mathcal{B}}_{E,g})\subseteq\overline{\mathcal{B}}_{E,g}$ denote the $p$-rank $\leq f$ locus of ¯$\overline{\mathcal{B}}_{E,g}$. To obtain a lower bound on $\dim V_{f}(\overline{\mathcal{B}}_{E,g})$ for $g\geq 2$, we use the purity of the $p$-rank stratification, a standard tool in this context. We rely on a version by Faber and van der Geer, based on a result of Oort, later generalized to the Newton polygon stratification as de Jong-Oort’s purity.
@@ 3.2 Computing the dimensions of the $p$-rank strata of ¯$\overline{\mathcal{B}}_{E,g}$
Let $B$ be an irreducible scheme over $k$ and let $\mathcal{D}\to B$ be a family of stable curves over $B$. Let $f$ be the $p$-rank of the generic fiber $D_{\eta}$. Then the closed subset of $B$ over which the $p$-rank of the fiber is at most $f-1$ is either empty or pure of codimension $1$ in $B$.
Let ¯$\overline{\mathcal{H}}_{g}$ denote the locus of stable hyperelliptic curves of genus $g$, and let $V_{f}(\overline{\mathcal{H}}_{g})\subseteq\overline{\mathcal{H}}_{g}$ be its $p$-rank $\leq f$ locus. Throughout the section, we will need the following result by Glass and Pries.
Lemma 3.5 ([GP05, Proposition 2.2], $p>2$).
Let $g\geq 2$. For any $0\leq f\leq g$, the locus $V_{f}(\overline{\mathcal{H}}_{g})$ is pure of dimension $g-1+f$ in ¯$\overline{\mathcal{H}}_{g}$.
@@
Let $E$ be an elliptic curve in characteristic $p>2$ with $p$-rank $f_{E}$. Recall that ¯${\overline{\mathcal{B}}}_{E,2}$ is irreducible and proper by Proposition 2.7 and Remark 3.2, while $\mathcal{B}_{E,2}$ cannot contain complete subvarieties by Remark 3.1 so that ¯$\overline{\mathcal{B}}_{E,2}-{\mathcal{B}}_{E,2}\neq\o$. One boundary point of ¯$\overline{\mathcal{B}}_{E,2}-{\mathcal{B}}_{E,2}$ corresponds to a curve $D^{\prime}=(E^{\prime},q_{1},q_{2})/\{q_{1}=q_{2}\}$ as in Example 3.3 with $f_{D^{\prime}}=f_{E}+1$. Therefore, by the purity result (Lemma 3.4), the generic $p$-rank of ¯${\overline{\mathcal{B}}}_{E,2}$ equals $f_{E}+1$.
It remains to show that $V_{f_{E}}(\overline{\mathcal{B}}_{E,2})$ is non-empty. By Example 3.3, $D=(E,O_{E})\cup_{O_{E}=O_{E}}(E,O_{E})$, with $f_{D}=2f_{E}$, represents a curve lying in the boundary of ¯${\overline{\mathcal{B}}}_{E,2}$. If $f_{E}=0$, then $f_{D}=0$. Otherwise, assume $f_{E}=1$. The locus of ordinary curves in ¯${\overline{\mathcal{B}}}_{E,2}$ is quasi-affine; this follows from [Oor99, Theorem 1.2] stating that the locally closed Ekedahl-Oort strata are quasi-affine applied to the final type $\nu=\{1,2\}$. Therefore, ¯${\overline{\mathcal{B}}}_{E,2}$ contains an element that corresponds to a curve $D$ with with lower $p$-rank, $f_{D}=f_{E}=1$. Moreover, $D$ is a smooth curve. This follows from Example 3.3, using that both $D_{1}$ and $D_{2}$ have $f_{D_{1}}=f_{D_{2}}=2$ when $f_{E}=1$. ∎
@@
It is enough to consider the case $f_{E}=0,f_{D}=0$ since the proof of the preceding lemma implies the result in other situations. Write $E:y^{2}=x(x-1)(x-\lambda)$. For $p>3$, there is $\lambda^{\prime}\neq\lambda$ such that $E^{\prime}:y^{2}=x(x-1)(x-\lambda^{\prime})$ is supersingular; see [Sil09, V.4]. The normalization of $E\times_{\mathbb{P}^{1}}E^{\prime}$ is a supersingular genus $2$ curve with a map of degree $2$ to $E$ and the result follows. ∎
@@
The first part of the proof of Lemma 3.6 establishes that there is a smooth curve $D$ of genus $2$ which is a double cover of an ordinary elliptic curve $E$ in characteristic $3$. However, if $E$ is a supersingular elliptic curve in characteristic $3$, the locus $V_{0}(\mathcal{B}_{E,2})$ of smooth curves $D$ in $\mathcal{B}_{E,2}$ with $f_{D}=0$ is empty. This follows from [Eke87, Theorem 1.1], which states that there are no smooth superspecial curves of genus $2$ in characteristic $3$ and the second part of the proof of Lemma 3.6.
@@
Lemma 3.10 ($p>2$).
For any $g\geq 2$ and any $2f_{E}\leq f\leq g-2+2f_{E}$ (resp. $f_{E}+1\leq f\leq g-1+f_{E}$), we have
@@
    $(E,O_{E})\cup_{O_{E}=q_{1}}(H,q_{1},q_{2})\cup_{q_{2}=O_{E}}(E,O_{E})$ representing an element of $V_{f}(\delta_{1,g-2})$, where $H$ represents an element of $V_{f-2f_{E}}(\overline{\mathcal{H}}_{g-2})$, and $q_{1},q_{2}\in H$ are two distinct points that are in the same orbit under the involution of $H$; or
@@
Let us fix $f_{E}\leq f\leq g-1+f_{E}$, $g_{1}\geq 2$ and $g_{2}\geq 1$ such that $g_{1}+g_{2}=g$. (Note that the outcomes of the computations below are also valid in the cases $g_{1}=g-1$ and $g_{1}=g-2$ since $\dim V_{0}(\overline{\mathcal{M}}_{0,3})=0$ and $\dim V_{f_{2}}(\overline{\mathcal{M}}_{1,1})=f_{2}$, and using the identifications ¯$\overline{\mathcal{H}}_{0,1}\cong\overline{\mathcal{M}}_{0,3}$ and ¯$\overline{\mathcal{H}}_{1}\cong\overline{\mathcal{M}}_{1,1}$ similarly as in Section 2.1 and the proof of Lemma 3.10.)
@@
and therefore, $\dim V_{f}(\Delta_{g_{1},g_{2}})=g-3+f-f_{E}$, for any $12\leq g_{1}\leq g-$ 1.
Now, we compute the dimension of $V_{f}(\Xi_{g_{1},g_{2}-1})$. Let $f_{1}$ and $f_{2}$ be numbers that satisfy $f_{E}\leq f_{1}\leq g_{1}$, $10\leq f_{2}\leq g_{2}-$ 1, and $f_{1}+f_{2}=f-1$. Let $D$ be a curve whose isomorphism class lies in $V_{f_{1}}(\overline{\mathcal{B}}_{E,g_{1}})$, let $H$ be a hyperelliptic curve from $V_{f_{2}}(\overline{\mathcal{H}}_{g_{2}-1})$, and $q_{1}\in D$ and $q_{1}^{\prime}\in H$ their points that are not ramification points of the structural morphisms $D\to E$ and $H\to\mathbb{P}^{1}$. Denote by $q_{2}\in D$ (resp. $q_{2}^{\prime}\in H$) the image of $q_{1}$ (resp. $q_{1}^{\prime}$) under the structural bielliptic (resp. hyperelliptic) involution. Every element of $V_{f}(\Xi_{g_{1},g_{2}-1})$ corresponds to a clutch of such curves $D$ and $H$ with $q_{1}$ and $q_{1}^{\prime}$ being identified, as well as $q_{2}$ and $q_{2}^{\prime}$. By the assumption we find that
@@
Theorem 3.12 ($p>2$).
@@
First, observe that $V_{f}(\overline{\mathcal{B}}_{E,g})$ is non-empty for any $f_{E}\leq f\leq g-1+f_{E}$. For $g=2$, this follows from Lemma 3.6, and for $g\geq 3$, it follows from Lemmas 3.6 and 3.5, which establish that $V_{f}(\Delta_{2,g-2})\subset V_{f}(\overline{\mathcal{B}}_{E,g})$ is non-empty. Following the strategy from Theorem [FvdG04, Theorem 2.3], let $W=W_{f}$ be a component of $V_{f}(\overline{\mathcal{M}}_{E,g})$ for a fixed $f$ with $f_{E}\leq f\leq g-1+f_{E}$. Let $W_{f+1}$ be a component of $V_{f+1}(\overline{\mathcal{B}}_{E,g})$ such that $W_{f}\subseteq W_{f+1}$. This gives us a sequence $W_{f}\subseteq W_{f+1}\subseteq\ldots\subseteq W_{g-2+f_{E}}\subseteq W_{g-1+f_{E}}=\overline{\mathcal{B}}_{E,g}$. Notice that $W_{g-1+f_{E}}=\overline{\mathcal{B}}_{E,g}$ since ¯$\overline{\mathcal{B}}_{E,g}$ is irreducible by Remark 3.2. By Lemma 3.4, we conclude that
@@
To establish an upper bound on the dimension of $W$, we use induction on $g$ with the base case $g=2$ covered by Lemma 3.6. The inductive hypothesis assumes that for all $2\leq g^{\prime}<g$ and $1\leq f^{\prime}\leq f$, $V_{f^{\prime}}(\overline{\mathcal{M}}_{E,g^{\prime}})$ is pure of dimension $g^{\prime}-2+f^{\prime}-f_{E}$. Since $\mathcal{B}_{E,g}$ cannot contain complete subvarieties (see Remark 3.1) it follows that $W\cap\Delta\neq\o$, where $\Delta=\Delta_{0}\cup\Delta_{1}\cup\ldots\cup\Delta_{\lfloor g/2\rfloor}$ is the boundary divisor of ¯$\overline{\mathcal{M}}_{g}$.
In the case $f_{E}=1$, we find that the underlying curve of the generic point of $W$ is smooth by comparing dimensions. Therefore, $W\cap\Delta_{i}\neq\o$ for some $0\leq i\leq\lfloor\frac{g}{2}\rfloor$. In Lemmas 3.10 and 3.11, we compute the dimensions of the loci of curves in ¯$\overline{\mathcal{B}}_{E,g}-\mathcal{B}_{E,g}$, under the assumption given by the inductive hypothesis, resulting in $\dim(W\cap\Delta_{i})\leq g-4+f$ for any $0\leq i\leq\lfloor\frac{g}{2}\rfloor$. Since $\Delta_{i}$ has codimension $1$ in ¯$\overline{\mathcal{M}}_{g}$ and ¯$\overline{\mathcal{M}}_{g}$ is a smooth stack, every component of $W\cap\Delta_{i}$ has codimension at most $1$ in $W$, and we get the following inequality:
@@
Consequently, we conclude that $\dim W=g-3+f=g-2+f-f_{E}$.

### v2

@@ 3.1 The boundary of $\mathcal{B}_{E,g}$
Let $\Delta=\overline{\mathcal{M}}_{g}-\mathcal{M}_{g}=\Delta_{0}\cup\Delta_{1}\cup\ldots\cup\Delta_{\lfloor g/2\rfloor}$, denote the boundary of $\overline{\mathcal{M}}_{g}$, with $\Delta_{i}$ its irreducible components. The generic point of $\Delta_{0}$ corresponds to an irreducible curve $C_{0}$ with an ordinary double point, whose normalization $\tilde{C}_{0}$ is a genus-(g1)(g-1) curve, while the generic point of $\Delta_{i}$ corresponds to a reducible curve with components of genus $i$ and gig-i for any $1\leq i\leq\left\lfloor\frac{g}{2}\right\rfloor$. By abuse of notation, we use $\Xi_{g_{1},g_{2}-1}$, $\Delta_{g_{1},g_{2}}$ for $g_{1},g_{2}\geq 0$ with $g_{1}+g_{2}=g$ to denote the image of $\Xi_{g_{1},g_{2}-1}$, $\Delta_{g_{1},g_{2}}$ under the forgetful morphism $\overline{M}_{E,g}^{{\rm red}}\to\overline{\mathcal{M}}_{g}$.
It follows from the description in Section 2.1 that each boundary component consists of isomorphism classes of curves that lie in the closure of a locus of double covers of the configuration
@@
with exactly 2g22g-2 distinct smooth branch points and that curves corresponding to elements in the loci $\Xi_{g_{1},g_{2}-1}$ (resp. $\Delta_{g_{1},g_{2}}$) are unramified (resp. ramified) above the node. We see that $\overline{\mathcal{B}}_{E,g}-\mathcal{B}_{E,g}$ consists of the union of the loci $\Xi_{g_{1},g_{2}-1}$ and $\Delta_{g_{1},g_{2}}$. The dimensions of these loci equal 2g42g-4 by Proposition 2.9. Let $\delta_{1,g-2}\subseteq\Xi_{1,g-2}$ denote the union of the components of $\Xi_{1,g-2}$ whose generic points are each represented by a curve of compact type, and denote $\xi_{1,g-2}=\Xi_{1,g-2}-\delta_{1,g-2}$. We have the following relations:
@@
Example 3.3 (Description of the boundary $\overline{\mathcal{B}}_{E,2}-\mathcal{B}_{E,2}$).
Let $E=(E,O_{E})$ be an arbitrary elliptic curve as above, and let $f_{E}$ be its $p$-rank. In Example 2.5, we describe the boundary of $\overline{M}_{E,2}-{M}_{E,2}$ of $M_{E,2}$, which we use to describe its image in $\overline{\mathcal{M}}_{2}$ here. The locus $\overline{\mathcal{B}}_{E,2}-\mathcal{B}_{E,2}$ is 0-dimensional and it consists of isomorphism classes of curves that are double covers of
@@
We note that there are exactly $3$ curves of the first kind and a unique curve of the second kind. Namely, the étale double covers of $E$ correspond to the choices of $2$-torsion points on $E$. Since #E[2](k)=4\#E[2](k)=4, there are $3$ non-trivial $2$-torsion points on $E$ defining different étale double covers $E^{\prime}\to E$, while $O_{E}$ is the fourth $2$-torsion point, and it defines a trivial double cover $(E,O_{E})\sqcup(E,O_{E})\to(E,O_{E})$. Since $f_{E^{\prime}}=f_{E}$, note that $f_{D_{1}}=f_{E}+1$ (see [BLR90, Section 9.2, Example 8]), while $f_{D_{2}}=2\cdot f_{E}$.
3.2 Computing the dimensions of the $p$-rank strata of $\overline{\mathcal{B}}_{E,g}$
Let $V_{f}(\overline{\mathcal{B}}_{E,g})\subseteq\overline{\mathcal{B}}_{E,g}$ denote the $p$-rank $\leq f$ locus of $\overline{\mathcal{B}}_{E,g}$. To obtain a lower bound on $\dim V_{f}(\overline{\mathcal{B}}_{E,g})$ for $g\geq 2$, we use the purity of the $p$-rank stratification, a standard tool in this context. We rely on a version by Faber and van der Geer, based on a result of Oort, later generalized to the Newton polygon stratification as de Jong-Oort’s purity.
@@ 3.2 Computing the dimensions of the $p$-rank strata of $\overline{\mathcal{B}}_{E,g}$
Let $B$ be an irreducible scheme over $k$ and let $\mathcal{D}\to B$ be a family of stable curves over $B$. Let $f$ be the $p$-rank of the generic fiber $D_{\eta}$. Then the closed subset of $B$ over which the $p$-rank of the fiber is at most f1f-1 is either empty or pure of codimension 11 in $B$.
Let $\overline{\mathcal{H}}_{g}$ denote the locus of stable hyperelliptic curves of genus $g$, and let $V_{f}(\overline{\mathcal{H}}_{g})\subseteq\overline{\mathcal{H}}_{g}$ be its $p$-rank $\leq f$ locus. Throughout the section, we will need the following result by Glass and Pries.
Lemma 3.5 ([GP05, Theorem 1, Proposition 2]).
Let $g\geq 2$. For any $0\leq f\leq g$, the locus $V_{f}(\overline{\mathcal{H}}_{g})$ is pure of dimension g1+fg-1+f in $\overline{\mathcal{H}}_{g}$.
@@
Let $E$ be an elliptic curve in characteristic $p>2$ with $p$-rank $f_{E}$. Recall that ${\overline{\mathcal{B}}}_{E,2}$ is irreducible and proper by Lemma 3.1, while $\mathcal{B}_{E,2}$ cannot contain complete subvarieties by Lemma 3.2. Therefore, $\overline{\mathcal{B}}_{E,2}-{\mathcal{B}}_{E,2}\neq\varnothing$. One boundary point of $\overline{\mathcal{B}}_{E,2}-{\mathcal{B}}_{E,2}$ corresponds to a curve $D^{\prime}=(E^{\prime},q_{1},q_{2})/\{q_{1}=q_{2}\}$} as in Example 3.3 with $f_{D^{\prime}}=f_{E}+1$ (computed using [BLR90, Section 9.2, Example 8]). Therefore, by the purity result (Lemma 3.4), the generic $p$-rank of ${\overline{\mathcal{B}}}_{E,2}$ equals $f_{E}+1$.
It remains to show that $V_{f_{E}}(\overline{\mathcal{B}}_{E,2})$ is non-empty. By Example 3.3, $D=(E,O_{E})\cup_{O_{E}=O_{E}}(E,O_{E})$, with $f_{D}=2f_{E}$, represents a curve lying in the boundary of ${\overline{\mathcal{B}}}_{E,2}$. If $f_{E}=0$, then $f_{D}=0$. Otherwise, assume $f_{E}=1$. The locus of ordinary curves in ${\overline{\mathcal{B}}}_{E,2}$ is quasi-affine; this follows from [Oor99, 6.1]. Therefore, ${\overline{\mathcal{B}}}_{E,2}$ contains an element that corresponds to a curve $D$ with lower $p$-rank, $f_{D}=f_{E}=1$. Moreover, DD is a smooth curve. This follows from Example 3.3, using that both $D_{1}$ and $D_{2}$ have $f_{D_{1}}=f_{D_{2}}=2$ when $f_{E}=1$. ∎
@@
It is enough to consider the case $f_{E}=0,f_{D}=0$ since the proof of the preceding lemma implies the result in other situations. Write $E:y^{2}=x(x-1)(x-\lambda)$. For $p>3$, there is $\lambda^{\prime}\neq\lambda$ such that $E^{\prime}:y^{2}=x(x-1)(x-\lambda^{\prime})$ is supersingular; see [Sil09, V.4]. Consider the double covers $E\to\mathbb{P}^{1}$ and $E^{\prime}\to\mathbb{P}^{1}$ given by the inclusion $k(x)\subset k(x,y)$ of function fields, and the fiber product $E\times_{\mathbb{P}^{1}}E^{\prime}$ taken with respect to these maps. The normalization of $E\times_{\mathbb{P}^{1}}E^{\prime}$ is a supersingular genus $2$ curve with a map of degree $2$ to $E$ and the result follows. ∎
@@
The first part of the proof of Lemma 3.6 establishes that there is a smooth curve $D$ of genus $2$ which is a double cover of an ordinary elliptic curve $E$ in characteristic $3$. However, if $E$ is a supersingular elliptic curve in characteristic $3$, the locus $V_{0}(\mathcal{B}_{E,2})$ of smooth curves $D$ in $\mathcal{B}_{E,2}$ with $f_{D}=0$ is empty. This follows from [Eke87, Theorem 1.1], which states that there are no smooth superspecial curves of genus $2$ in characteristic $3$ and Remark 3.7.
@@
Lemma 3.10.
For any $g\geq 2$ and $2f_{E}\leq f\leq g-2+2f_{E}$ (resp. $f_{E}+1\leq f\leq g-1+f_{E}$), we have
@@
    $(E,O_{E})\cup_{O_{E}=q_{1}}(H,q_{1},q_{2})\cup_{q_{2}=O_{E}}(E,O_{E})$ representing an element of $V_{f}(\delta_{1,g-2})$, where $H$ represents an element of $V_{f-2f_{E}}(\overline{\mathcal{H}}_{g-2})$, and $q_{1},q_{2}\in H$ are two distinct points that are in the same orbit under the involution of $H$.
@@
Let us fix $f_{E}\leq f\leq g-1+f_{E}$, $g_{1}\geq 2$ and $g_{2}\geq 1$ such that $g_{1}+g_{2}=g$. (Note that the outcomes of the computations below are also valid in the cases $g_{1}=g-1$ and $g_{1}=g-2$ since $\dim V_{0}(\overline{\mathcal{M}}_{0,3})=0$ and $\dim V_{f_{2}}(\overline{\mathcal{M}}_{1,1})=f_{2}$, and using the identifications $\overline{\mathcal{H}}_{0,1}\cong\overline{\mathcal{M}}_{0,3}$ and $\overline{\mathcal{H}}_{1}\cong\overline{\mathcal{M}}_{1,1}$ similarly as in Section 2.1 and the proof of Lemma 3.10.)
@@
and therefore, $\dim V_{f}(\Delta_{g_{1},g_{2}})=g-3+f-f_{E}$, for any $2\leq g_{1}\leq g-1$.
Now, we compute the dimension of $V_{f}(\Xi_{g_{1},g_{2}-1})$. Let $f_{1}$ and $f_{2}$ be numbers that satisfy $f_{E}\leq f_{1}\leq g_{1}$, $0\leq f_{2}\leq g_{2}-1$, and $f_{1}+f_{2}=f-1$. Let $D$ be a curve whose isomorphism class lies in $V_{f_{1}}(\overline{\mathcal{B}}_{E,g_{1}})$, let $H$ be a hyperelliptic curve from $V_{f_{2}}(\overline{\mathcal{H}}_{g_{2}-1})$, and $q_{1}\in D$ and $q_{1}^{\prime}\in H$ their points that are not ramification points of the structural morphisms $D\to E$ and $H\to\mathbb{P}^{1}$. Denote by $q_{2}\in D$ (resp. $q_{2}^{\prime}\in H$) the image of $q_{1}$ (resp. $q_{1}^{\prime}$) under the structural bielliptic (resp. hyperelliptic) involution. Every element of $V_{f}(\Xi_{g_{1},g_{2}-1})$ corresponds to a clutch of such curves $D$ and $H$ with $q_{1}$ and $q_{1}^{\prime}$ being identified, as well as $q_{2}$ and $q_{2}^{\prime}$. By the assumption we find that
@@
Before proving our main result, let us introduce some notation we will use in the proof of the case $f_{E}=0$. Let $j:\overline{\mathcal{M}}_{g}\to\tilde{\mathcal{A}}_{g}$ denote the Torelli morphism, where $\tilde{\mathcal{A}}_{g}$ is a fixed smooth toroidal compactification of the moduli space of principally polarized abelian varieties $\mathcal{A}_{g}$, as in [Ale04]. We can now turn to the proof of our main result.
Theorem 3.12.
@@
First, observe that $V_{f}(\overline{\mathcal{B}}_{E,g})$ is non-empty for any $f_{E}\leq f\leq g-1+f_{E}$. For $g=2$, this follows from Lemma 3.6, and for $g\geq 3$, it follows from Lemmas 3.5 and 3.6, which establish that $V_{f}(\Delta_{2,g-2})\subset V_{f}(\overline{\mathcal{B}}_{E,g})$ is non-empty.
We follow the strategy of the proof of Theorem [FvdG04, Theorem 2.3]. Let $W=W_{f}$ be a component of the $p$-rank $\leq f$ locus $V_{f}(\overline{\mathcal{B}}_{E,g})$ for a fixed $f$ with $f_{E}\leq f\leq g-1+f_{E}$. Let $W_{f+1}$ be a component of $V_{f+1}(\overline{\mathcal{B}}_{E,g})$ such that $W_{f}\subseteq W_{f+1}$. This gives us a sequence
$W_{f}\subseteq W_{f+1}\subseteq\ldots\subseteq W_{g-2+f_{E}}\subseteq W_{g-1+f_{E}}.$
Note that $W_{g-1+f_{E}}=\overline{\mathcal{B}}_{E,g}$ and $\dim W_{g-1+f_{E}}=2g-3$ by Lemma 3.1. By Lemma 3.4, if $W_{i}\neq W_{i+1}$, for $f\leq i\leq g-2+f_{E}$, then $W_{i}$ has codimension 11 in $W_{i+1}$. This implies that the codimension of $W=W_{f}$ in $\overline{\mathcal{B}}_{E,g}=W_{g-1+f_{E}}$ is at most $g-1+f_{E}-f$, or equivalently, that
@@
To establish an upper bound on the dimension of $W$, we use induction on $g$ with the base case $g=2$ covered by Lemma 3.6. The inductive hypothesis assumes that for all $2\leq g^{\prime}<g$ and $1\leq f^{\prime}\leq f$, $V_{f^{\prime}}(\overline{\mathcal{M}}_{E,g^{\prime}})$ is pure of dimension $g^{\prime}-2+f^{\prime}-f_{E}$. Since $\mathcal{B}_{E,g}$ cannot contain complete subvarieties (see Lemma 3.2) it follows that $W\cap\Delta\neq\varnothing$, where $\Delta=\Delta_{0}\cup\Delta_{1}\cup\ldots\cup\Delta_{\lfloor g/2\rfloor}$ is the boundary divisor of $\overline{\mathcal{M}}_{g}$.
In the case $f_{E}=1$, we find that the underlying curve of the generic point of $W$ is smooth by comparing dimensions. Therefore, $W\cap\Delta_{i}\neq\varnothing$ for some $0\leq i\leq\lfloor\frac{g}{2}\rfloor$. In Lemmas 3.10 and 3.11, we compute the dimensions of the loci of curves in $\overline{\mathcal{B}}_{E,g}-\mathcal{B}_{E,g}$, under the assumption given by the inductive hypothesis, resulting in $\dim(W\cap\Delta_{i})\leq g-4+f$ for any $0\leq i\leq\lfloor\frac{g}{2}\rfloor$. Since $\Delta_{i}$ has codimension 11 in $\overline{\mathcal{M}}_{g}$ and $\overline{\mathcal{M}}_{g}$ is a smooth stack, every component of $W\cap\Delta_{i}$ has codimension at most 11 in $W$ (see [Vis89, page 614]), and we get the following inequality:
@@
Consequently, we conclude that $\dim W\leq g-3+f=g-2+f-f_{E}$.